Solve the system of equations and give the value of x:
3x + 12y = 9
x = -4y - 3
Please show details thank you so much.

Answer by oscargut(2103)

3x + 12y = 9
x = -4y - 3
Then:
3x + 12y = 9
x +4y = -3
If we multiply the 2nd equation by -3 we have :
3x + 12y = 9
-3x -12y = 9
Adding them:
0x+0y = 18 (not true)
Answer: The system has no solution

If you need more help with math questions/assignments/tests
You can ask me more using this e-mail (mthman@gmail.com) or you can visit my site:
http://mathshelp-online.com/
(Payment is using paypal)